OpenStudy (jewlzme17):

A cylinder has a height that is 2 times as large as its radius. The lateral area of the cylinder is 16 {pi} square units. What is the length of the radius of the cylinder?

OpenStudy (superdavesuper):

The lateral area = diameter x height okay? so what will that tell u?

OpenStudy (superdavesuper):

Oooops....sorry wrong equation! it should be lateral area = pi x diameter x height....

OpenStudy (anonymous):

Lateral area \[A(r)=2 \pi r h = 4 \pi r^2=16 \pi \implies r=2\]is

OpenStudy (jewlzme17):

so s that the answer?

OpenStudy (jewlzme17):

@AnimalAin

OpenStudy (anonymous):

Let radius = r units ==> ht = 2r units A = 2πrh 2πr(2r) = 16π 4r^2 = 16 r = 4 units

OpenStudy (superdavesuper):

AnimalAin has the right ans